GNU autosprintf, version 1.0

Formatted Output to Strings in C++

Bruno Haible

Copyright (C) 2002-2003, 2006-2007 Free Software Foundation, Inc. This manual is free documentation. It is dually licensed under the GNU FDL and the GNU GPL. This means that you can redistribute this manual under either of these two licenses, at your choice. This manual is covered by the GNU FDL. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License (FDL), either version 1.2 of the License, or (at your option) any later version published by the Free Software Foundation (FSF); with no Invariant Sections, with no Front-Cover Text, and with no Back-Cover Texts. A copy of the license is at http://www.gnu.org/licenses/fdl.

html.

This manual is covered by the GNU GPL. You can redistribute it and/or modify it under the terms of the GNU General Public License (GPL), either version 2 of the License, or (at your option) any later version published by the Free Software Foundation (FSF). A copy of

the license is at http://www.gnu.org/licenses/gpl.html.

1 Introduction

This package makes the C formatted output routines (fprintf et al.) usable in C++ programs, for use with the <string> strings and the <iostream> streams.

It allows to write code like

      cerr << autosprintf ("syntax error in %s:%d: %s", filename, line, errstring);

instead of

      cerr << "syntax error in " << filename << ":" << line << ": " << errstring;

The benefits of the autosprintf syntax are:

• It reuses the standard POSIX printf facility. Easy migration from C to C++.

• English sentences are kept together.

• It makes internationalization possible. Internationalization requires format strings, because in some cases the translator needs to change the order of a sentence, and more generally it is easier for the translator to work with a single string for a sentence than with multiple string pieces.

• It reduces the risk of programming errors due to forgotten state in the output stream

    (e.g. cout << hex; not followed by cout << dec;).

2 The autosprintf class

An instance of class autosprintf just contains a string with the formatted output result. Such an instance is usually allocated as an automatic storage variable, i.e. on the stack, not with new on the heap.

The constructor autosprintf (const char *format, ...) takes a format string and additional arguments, like the C function printf.

Conversions to char * and std::string are defined that return the encapsulated string. The conversion to char * returns a freshly allocated copy of the encapsulated string; it needs to be freed using delete[]. The conversion to std::string returns a copy of the encapsulated string, with automatic memory management.

The destructor ~autosprintf () destroys the encapsulated string.

An operator << is provided that outputs the encapsulated string to the given ostream.

3 Using autosprintf in own programs

To use the autosprintf class in your programs, you need to add

      #include "autosprintf.h"
      using gnu::autosprintf;

to your source code. The include file defines the class autosprintf, in a namespace called gnu. The ‘using’ statement makes it possible to use the class without the (otherwise natural) gnu:: prefix.

When linking your program, you need to link with libasprintf, because that’s where the class is defined. In projects using GNU autoconf, this means adding

‘AC_LIB_LINKFLAGS([asprintf])’ to configure.in or configure.ac, and using the

@LIBASPRINTF@ Makefile variable that it provides.